Write 116 2/3 as a fraction or a mixed number.
help someone please

Found 2 solutions by checkley71, edjones:
Answer by checkley71(8403)

116 2/3=(3*116+2)/3=(348+2)/3=350/3 answer.

Answer by edjones(8007)

It is a mixed number already.
116 2/3
3*116+2=350 Multiply the denominator by the whole number and add the numerator.
Memorize how to do this. you will use it the rest of you life.
350/3 Now it is a fraction.
.
Ed